# New entropy, thermodynamics of apparent horizon and cosmology

### v1

@@ IV Generalized Friedmann’s equations
Figure 7 shows that when $z$ increases $q$ also increases. When EoS parameter for the matter $w$ increases, at fixed $b$ (Left panel), the deceleration parameter $q$ also increases. At $w=-0.3$ there are two phases, universe acceleration $q<0$ and deceleration $q>0$ but at $w=0$ and $w=1/3$ one has only the deceleration phase. According to Fig. 5 (Right panel), when parameter $b$ increases at fixed $w$ the deceleration parameter $q$ decreases. One has here two phases, acceleration and deceleration for $b=0.4$ and $b=0.6$ but for $b=0.2$ we have the deceleration phase.
V Conclusion
@@
Considering a novel entropy $S_{K}=S_{BH}/(1+\gamma S_{BH})$ we have a property similar to the Bekenstein–Hawking entropy $S_{BH}$. By using our entropy and by employing the holographic principle, we have obtained a dark energy model. The $S_{K}$ vanishes when the apparent horizon radius $R_{h}$ is zero and it increases monotonically when the apparent horizon radius $R_{h}$ increases. The barotropic perfect fluid and flat FLRW universe were studied. We have obtained, from first law of apparent horizon thermodynamics, the modified Friedmann’s equations. The second Friedmann’s equation possesses an addition term corresponding to the density of dark energy which can be treated as a dynamical cosmological constant. Implying that there is no interaction between various components of cosmos, and the dark energy density and pressure $\rho_{D}$ and $p_{D}$ obeys ordinary conservation law, we computed $\rho_{D}$ and $p_{D}$. The EoS parameter for dark energy $w_{D}=p_{D}/\rho_{D}$ was calculated and it was shown that $\lim_{H\rightarrow\infty}w_{D}=-1$. The universe can have two phases, acceleration and deceleration, that is due to holographic dark energy. It should be noted that Barrow and Tsallis entropies also lead to cosmology which ia due to Einsten’s equations with the dynamical cosmological constant Gennaro . We computed the deceleration parameter that shows the possibility, for some model parameters, to describe inflation and acceleration at current era. The transition redshifts (when $q=0$) was calculated (given in Table I) for some parameters. We show (see Table II) that at $b\approx 0.5$ and $w=-2/3$ we have the deceleration parameter $q_{0}\approx-0.6$ which was observed in universe at the current era Roos . Thus, our approach, based on new entropy and leading to modified Friedmann equations, can description universe inflation and late time universe acceleration. It is worth mentioning that the results obtained here are similar to those in another model with different entropy Kruglov .
Data availability statement: No data was used for the research described in the article.

### v3

@@ 4 Generalized Friedmann’s equations
Figure 6 shows that when $z$ increases $q$ also increases. When EoS parameter for the matter $w$ increases, at fixed $b$ (Left panel), the deceleration parameter $q$ also increases. At $w=-0.3$ there are two phases, universe acceleration $q<0$ and deceleration $q>0$ but at $w=0$ and $w=1/3$ one has only the deceleration phase. According to Fig. 6 (Right panel), when parameter $b$ increases at fixed $w$ and $z$, the deceleration parameter $q$ decreases. One has here two phases, acceleration and deceleration for $b=0.4$ and $b=0.6$ but for $b=0.2$ we have the deceleration phase.
4.1 Inflation
Inflation corresponds to the accelerated expansion of the universe and gravity possesses a repulsive force, At the early stage of the universe, i.e. inflation, the matter is absent, $\rho=p=0$. The entropic energy (dark energy) density triggers the universe to accelerate. Then we obtain from Eq. (16) that $\dot{H}=0$, i.e. during inflation the Hubble parameter is a constant. As a result, the scale factor is given by $a(t)=a_{0}\exp(Ht)$ and corresponds to de Sitter spacetime. It should be noted that de Sitter stage corresponds to the eternal inflation. The Hubble parameter at the early universe (inflation) is $H\approx 10^{-3}M_{Pl}$ ($M_{Pl}=1/\sqrt{16\pi G}$). Then $GH^{2}\approx 2\cdot 10^{-8}$. From Eq. (17), at $\rho=0$ we obtain the equation as follows:
$H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=0.$ (32)
Introducing new variable $x=H^{2}/b$, Eq. (32) becomes
$x-\frac{1}{x+1}-2\ln(x+1)=0.$ (33)
The non-trivial solution to Eq. (33) is given by $x=3.04002$ and we have the relation
$H^{2}=3.04002b.$ (34)
Knowing the Hubble parameter during universe inflation, $H\approx 10^{-3}M_{Pl}$, we can fix parameter $b$ from Eq. (34) and the entropy parameter $\gamma=bG/\pi$. Thus, we obtain $\gamma\approx 2\times 10^{-9}$. It is worth noting that to have a viable model of inflation one needs a smooth graceful exit into the deceleration stage [1].
5 F(T)-gravity from generalized entropy
In teleparallel gravity, a scalar torsion $T$ plays a role of a fundamental field similar to the curvature $R$ in Einstein’s general relativity theory [42, 43]. To describe the inflationary era and the current universe accelerating expansion, one introduces a Lagrangian in the form of $F(T)$ in analogy with $F(R)$-gravity. The torsion scalar field $T$ is given by [44, 45]
$T=S_{\rho}^{~{}\mu\nu}T^{\rho}_{~{}\mu\nu},$ (35)
where tensors $S_{\rho}^{~{}\mu\nu}$ and $T^{\rho}_{~{}\mu\nu}$ are defined as follows:
$S_{\rho}^{~{}\mu\nu}=\frac{1}{2}\left(K_{~{}~{}\rho}^{\mu\nu}+\delta^{\mu}_{\rho}T_{~{}~{}\alpha}^{\alpha\nu}-\delta^{\nu}_{\rho}T_{~{}~{}\alpha}^{\alpha\mu}\right),$
$K^{\mu\nu}_{~{}~{}\rho}=-\frac{1}{2}\left(T^{\mu\nu}_{~{}~{}\rho}-T^{\nu\mu}_{~{}~{}\rho}-T^{~{}\mu\nu}_{\rho}\right),$
$T^{\rho}_{~{}\mu\nu}=e^{\rho}_{i}\left(\partial_{\mu}e^{i}_{\nu}-\partial_{\nu}e^{i}_{\mu}\right),$ (36)
where $e^{i}_{\nu}$ ($i=0,1,2,3$) is a vierbein field, so that the metric tensor is $g_{\mu\nu}=\eta_{ij}e^{i}_{\mu}e^{j}_{\nu}$ with $\eta_{ij}$ being the flat metric of the tangent spacetime. For FLRW metric (5), we have $e^{i}_{\mu}=\mbox{diag}(1,a,a,a)$. Then the torsion scalar field is given by $T=-6H^{2}$. From the action by variation with respect to $e^{i}_{\mu}$, where the Lagrangian is $F(T)$, one obtains [46]
$\frac{1}{6}\left[F(T)-2TF^{\prime}(T)\right]|_{T=-6H^{2}}=\left(\frac{8\pi G}{3}\right)\rho.$ (37)
Making use of equations (17) and (37) we find
$\left[F(T)-2TF^{\prime}(T)\right]=-T-\frac{36b^{2}}{6b-T}-12b\ln\left(1-\frac{T^{2}}{6b}\right).$ (38)
By integration of Eq. (38) we obtain the function $F(T)$:
$F(T)=T-3\sqrt{6bT}~{}\mbox{tanh}^{-1}\sqrt{\frac{T}{6b}}$
$-12b\ln\left(1-\frac{T}{6b}\right)+C\sqrt{T}-6b,$ (39)
where $C$ is the integration constant. Thus, the teleparallel gravity with the function (39) corresponds to entropic cosmology with entropy (15) under consideration.
6 Conclusion
@@
Considering a novel entropy $S_{K}=S_{BH}/(1+\gamma S_{BH})$ we have a property similar to the Bekenstein–Hawking entropy $S_{BH}$. By using our entropy and by employing the holographic principle, we have obtained a dark energy model. The $S_{K}$ vanishes when the apparent horizon radius $R_{h}$ is zero and it increases monotonically when the apparent horizon radius $R_{h}$ increases. The barotropic perfect fluid and spatial flat FLRW universe were studied. We have obtained, from first law of apparent horizon thermodynamics, the modified Friedmann’s equations. The second Friedmann’s equation possesses an addition term corresponding to the density of dark energy which can be treated as a dynamical cosmological constant. Implying that there is no interaction between various components of cosmos, and the dark energy density and pressure $\rho_{D}$ and $p_{D}$ obeys ordinary conservation law, we computed $\rho_{D}$ and $p_{D}$. The EoS parameter for dark energy $w_{D}=p_{D}/\rho_{D}$ was calculated and it was shown that $\lim_{H\rightarrow\infty}w_{D}=-1$. Thus, at the small apparent horizon radius $R_{h}$ the de Sitter stage is realized which describes the inflation. In the model under consideration the universe can have two phases, acceleration and deceleration, that is due to holographic dark energy. It should be noted that Barrow and Tsallis entropies also lead to cosmology which ia due to Einstein’s equations with the dynamical cosmological constant [40]. We computed the deceleration parameter that shows the possibility, for some model parameters, to describe the acceleration at the current era. The transition redshifts (when $q=0$) was calculated (given in Table I) for some parameters. We show (see Table II) that at $b\approx 0.1$ and $w=-2/3$ we have the deceleration parameter $q_{0}\approx-0.535$ which was observed in universe at the current era [39]. Thus, our approach, based on new entropy and leading to modified Friedmann equations, can describe universe inflation and late time universe acceleration. It has been proven that entropic cosmology with our entropy proposed is equivalent to cosmology based on the teleparallel gravity with the function $F(T)$.